Reports 1-1 of 1 Clear search Modify search
Injection system (General activities)
derossi, pillant - 14:15 Tuesday 11 June 2019 (46075) Print this report
INJ activities during the maintenance

this morning during the maintenance we performed the following activities:

- IMC TF measurements (when the IMC reflection is used as the error signal). The TF looks good, the UGF is around 130 kHz and it is the same before and after performing other activities (BPC realignment in particular). It is reported in the 1st and 2nd plots.

- SL scan. We acquired again the signals to monitor the piezo on the SL ramepauto (the cables were unplugged, they are named PSL_SL_PZT_LF and PSL_SL_PZT_HF), and  we performed a scan at around 08:22:29 UTC (3rd plot)

- BPC corrections reduction (see 4th plot)

- we reduced the temperature of the laser amplifier diodes (plot 5). They had increased during these last weeks and we were close to the maximum. This is probably due to some impurities in the water circulating to cool the diodes. We just quickly unplugged and plugged again the cooling circuit from the chiller and the temperature started to drop down (probably modifying a bit the flux we made the impurities move out from the filter and now the water is better circulating)

Images attached to this report
Search
×

Warning

×